OpenStudy (anonymous):

Hi, could someone help me with this? (Please reply so i can draw some triangles that go along with the problem). Thanks.

OpenStudy (anonymous):

yes go on

OpenStudy (anonymous):

thanks, im drawing.

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

I have to draw first, but.. Triangle RST and JKL are similar. Which choice shows the equations that can be used to find the area of triangle JKL?

OpenStudy (anonymous):

A. First use 9/18=h/12 and then use area= 1/2 (12h)

OpenStudy (anonymous):

do width times hieght and figure out the height multiply it by the width on both triangles and compare

OpenStudy (anonymous):

B. First use 9/18 = h/12 and then use area= 12h

OpenStudy (anonymous):

Wait please, I am going to draw.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well, there's these 2 triangles. RST has a height of 9 cm and a base of 18 cm and JKL has a height of h cmand base of 12 cm.

OpenStudy (anonymous):

then do 9 times 18 and get ur answer add that togeter with 12 and a triangle shld be 180 degrees so then do 180 divided by 12 and u get ur answer